OpenStudy (anonymous):

The graph of f(x) = 2x + 4 shifts five units to the right when it is replaced with the graph of f(x) = 2x - k. What is the value of k? -5 -1 1 5 @keke_luvs_u

OpenStudy (anonymous):

@josh543 The answer is 6. but its not an option. Im positive that im right so,you should go with 5 sense it comes before 6?

OpenStudy (anonymous):

k

OpenStudy (jdoe0001):

\(\large { \textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Up\\ {\color{green}{ D}} < 0 & Down\end{array} \\ \qquad \downarrow\qquad\qquad\quad\ \downarrow\\ y = {\color{purple}{ A}} ( {\color{blue}{ 2}}x + {\color{red}{ 4}} ) + {\color{green}{ D}}\\ \qquad\qquad\quad\ \uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to \ left\\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to \ right\end{array} \end{array} }\) any ideas? bear in mind that the "parent" function is y = 2x + 4